OpenStudy (anonymous):

find the median -9,-3.1,4,5,10.2

OpenStudy (turingtest):

-9 and 10.2 are the extrema, so 10.2-4=6.2 9-4=5 10.2-3.1=7.1 9-3.1=6.9 looks like 4 is the median

OpenStudy (amistre64):

4 it is; but i got no idea what that other stuff is :)

OpenStudy (turingtest):

wait, do I remember the definition of "median" I was trying to minimize the distance from the extrema by trying the difference in the absolute value of the extrema minus 4 or -3.1, the two likely candidates

OpenStudy (amistre64):

median, the grassy spot in the middle of the road

OpenStudy (turingtest):

hardy har... :)

OpenStudy (anonymous):

yes. 4 is the median because you have listed the numbers from lowest to highest. Next, choose the number that is in the middle. 4 is your answer. For additional help, visit my website at http://tmtutoringservices.com